A sphere of radius 6cm is moulded into a think cylindrical wire of length 32cm. Calculate the radius of the wire in SI units

1
Expert's answer
2022-01-02T18:24:22-0500

"r=\sqrt{4/3\cdot π R^3/L}=0.005\text{ m}."